\begin{thebibliography}{}

\bibitem{AHT}
S. Adly, T. Haddad and L. Thibault,
 Convex sweeping process in the framework of measure differential
  inclusions and evolution variational inequalities.
 Math. Program, 148 (1-2, Ser. B) (2014) 5-47.

\bibitem{AL} S. Adly and    B. K. Le,  Unbounded second-order state-dependent Moreau's sweeping processes in Hilbert
spaces. J. Optim. Theory Appl. 169, (2016)  407-423.

\bibitem{AL1} S. Adly and    B. K. Le, Unbounded state-dependent sweeping
process with perturbations in uniformly convex and q-uniformly
smooth Banach spaces. Numer. Algebra Control Optim. 8, 81–95 (2018).

\bibitem{AN}  S. Adly and   F. Nacry,  An existence result for
discontinuous second-order nonconvex state-dependent sweeping
process. Appl. Math. Optim.  DOI 10.1007/s00245-017-9446-9 (2017).

 \bibitem{ANT} S. Adly, F. Nacry and L. Thibault, Discontinuous sweeping process with prox-regular sets.
 ESAIM: Control, Optimizatios and Calculus of Variation. 23 (4)
 (2017) 1293-1329.

 \bibitem{AA1}F. Aliouane, D. Azzam-Laouir, A second order differential
inclusion with proximal normal cone in Banach spaces, Topol Methods
Nonlinear Anal. 44(1) 143-160 (2014).


\bibitem{AA2}F. Aliouane, D. Azzam-Laouir, Second-order sweeping process with
a lipschitz perturbation. J. Math. Anal. Appl. 452, 729-746 (2017).

\bibitem{AACM}F. Aliouane, D. Azzam-Laouir, C. Castaing and M.D.P. Monteiro
Marques, Second-order time and state sweeping process in Hilbert
space, J Optimiz Theory App. 182, 153-188 (2019).

\bibitem{Az} D. Azzam-Laouir, A measure differential inclusion with time-dependent maximal monotone
 operators. ArXiv preprint,  arXiv:2103.07376 (2021) 1-15.

\bibitem{Az1} D. Azzam-Laouir, Mixed semicontinuous perturbation of a
second order nonconvex sweeping process. Electronic Journal of
Qualitative Theory of Differential Equations, 37, 1-9 (2008).


\bibitem{AI} D. Azzam-Laouir and S. Izza, Existence of solutions for second-order
perturbed nonconvex sweeping process. Computers \& Mathematics with
Applications. 62, 1736-1744 (2011).



\bibitem{ABCM1} D. Azzam-Laouir, W. Belhoula, C. Castaing and M.D.P. Monteiro Marques,
 Perturbed evolution problems with absolutely continuous variation in
time and applications. J. Fixed Point Theory Appl. (2019) 21: 40.


\bibitem{ABCM2} D. Azzam-Laouir, W. Belhoula, C. Castaing and M.D.P. Monteiro Marques,
Multi-valued perturbation to evolution problems involving time
dependent maximal monotone operators. Evolution Equations \& Control
Theory, 9(1) (2020) 219-254.

\bibitem{AB} D. Azzam-Laouir, I. Boutana-Harid,
        Mixed semicontinuous perturbation to an evolution problem with
        time-dependent maximal monotone operator. J.  Nonlinear  and Convex Analysis. Vol 20, Number 1 (2018)  39-52.


\bibitem{ACM} D. Azzam-Laouir, C. Castaing and M.D.P. Monteiro Marques,
Perturbed evolution proplems with continuous bounded variation in
time and applications. Set-Valued Var. Anal. Volume 26, No. 3 (2018)
693-728.


\bibitem{ACM1} D. Azzam-Laouir, C. Castaing and M.D.P. Monteiro
Marques, BV right-continuous solutions of differential inclusions
involving  time dependent maximal monotone operators. ArXiv:
2103.01113v1 (2021) 1-43.

\bibitem{Ba} V. Barbu,   Nonlinear Semigroups
and Differential Equations in Banach Spaces. Noordhoff Int. Publ.
Leyden,  (1976).

\bibitem{BCG} H. Benabdellah, C. Castaing and M.A. Gamal Ibrahim, BV Solutions of multivalued differential
equations on closed moving sets in Banach spaces. Banach center
publications, Vol 32, Institute of Mathematics, Polish academy of
Sciences, Warszawa (1995).


\bibitem{BCSS} H. Benabdellah,  C. Castaing, A. Salvadori and A. Syam, Nonconvex sweeping process.
 Journal of Applied  Analysis, Vol 2, No2 (1996) 17-40.


\bibitem{BAC} M. Benguessoum, D. Azzam-Laouir and C. Castaing,
On a  time and state dependent maximal monotone operator coupled
with a sweeping process with perturbations.  Set-Valued Var. Anal,
29 (2021) 191-219.

\bibitem{Bou} M. Bounkhel, Existence results for second order convex sweeping
processes in p-uniformly smooth and q-uniformly convex Banach
spaces, Electron J Qual Theo. 27, 1-10 (2012).

\bibitem{BL}
M. Bounkhel, D. Laouir-Azzam, Existence results for second-order
nonconvex sweeping processes. Set-Valued Var Anal. 12(3), 291-318
(2004).

\bibitem{brezis} H. Brezis, Op\'erateurs maximaux monotones, North Holland
    Publ. Compagny, Amsterdam-London, (1973).


    \bibitem{CIT} Camlibel K, Ianelli L, Tanwani A. Convergence of
proximal solutions for evolution inclusions with time-dependent
maximal monotone operators. Math. Programming, Series A.
https://doi.org/10.1007/s10107-021-01666-7 (2021)





    %\bibitem{CMA2} C. Castaing and M. D. P. Monteiro  Marques, BV Periodic solutions of an evolution problem associated with continuous
%convex sets. Set-Valued Anal. Vol 3 (1995) 381-399.

\bibitem{C1}C. Castaing, Quelques probl\`{e}mes d'\'{e}volution du
second ordre. Sem. Anal. Convexe. Montpellier. Expose No. 5. (1988).

\bibitem{CIY} C. Castaing, A.G. Ibrahim, M.F. Yarou, Some
contributions to nonconvex sweeping process. J. Nonlinear Convex
Anal. Vol. 10, No. 1, 1-20 (2009).

\bibitem{CM} C. Castaing, M.D.P. Monteiro Marques, Evolution problems
associated with nonconvex closed moving sets with bounded variation.
Portugal. Math. 53   (1996) 73-87.


\bibitem{CMRF} C. Castaing, M. D. P. Monteiro  Marques and P. Raynaud~de Fitte, Second order evolution
problems with time dependent maximal monotone  operator. Adv. Math.
Econ.  Vol 23  (2018).


%\bibitem{CV} C. Castaing and M. Valadier,  Convex Analysis and Measurable
%Multifunctions. Lectures Notes in Mathematics, Springer-Verlag,
%Berlin,  580 (1977).






    \bibitem{ET1}  J.F. Edmond and L. Thibault, BV solutions of  non convex sweeping proccess
differential inclusion with perturbation. J. Differential Equations.
22 (2006) 135-179.


    \bibitem {Ken}  N. Kenmochi, Solvability of nonlinear evolution equations with time-dependent constraints
 and applications. Bull. Fac. Educ. Chiba Univ. 30 (1981).






    \bibitem{KM} M. Kunze and M.D.P. Monteiro Marques, BV solutions to
evolution problems with time-dependent domains. Set-Valued Anal.
 5 (1997) 57-72.


 \bibitem{Le} B.K. Le, Well-posedness and nonsmooth Lyapunov pairs
 for state-dependent maximal monotone differential inclusions.
 Optimization, 69 (6) (2020) 1187-1217.


 \bibitem{Mon}
M.D.P. Monteiro Marques, Differential inclusions nonsmooth
mechanical problems, shocks and dry friction. Progress in Nonlinear
Differential Equations and Their Applications, Birkhauser, Vol 9
(1993).

\bibitem{M} J.J. Moreau, Rafle par un convexe variable I, Sem. Anal. Convexe
Montpellier (1971) Expos\'e 15.



\bibitem{M1} J.J. Moreau, Evolution
 problem asssociated with a moving convex set in a Hilbert space.
 J. Differiential. Equations. Vol 26  (1977) 347-374.

\bibitem{M2} J.J. Moreau, Bounded variation in time, in Topics in Non-smooth
Mechanics (J.J. Moreau, P.D. Panagiotopoulos, G. Strang, Eds.),
Birkh\"{a}user Verlag, Basel-Boston-Berlin (1988).


 \bibitem{M3} J.J. Moreau, Sur les mesures differentielles des fonctions
vectorielles \`a variation born\'ee. Sem. Anal. Convexe. Montpellier
  (1975) Expos\'e 17.


\bibitem{MV} J.J. Moreau and M. Valadier, A chain rule involving vector functions of bounded variations. J. Functional Analysis,
Vol 74, 2 (1987) 333-345.


\bibitem{NNT}  F. Nacry, J. Noel and L. Thibault,   On first and second order state-dependent prox-regular sweeping process.
 Preprint (2020).

 \bibitem{N} J. Noel,  Second-order general perturbed sweeping process
differential inclusion. J. Fixed Point Theory Appl. 20, 133 (2018).

 \bibitem{P} N.H. Pavel, Nonlinear
evolution operators and semigroups, Lecture Notes in Mathematics,
1260. Springer, Nex York (1987).


 \bibitem{Por} J.E. Porter, Helly selection principle for functions of bounded
$P$-variation. Rocky Mountain Journal of Mathematics, Vol 25, Number
2 (2005).


\bibitem{Saidi1} S. Saidi, On a second-order functional evolution problem with time and state dependent maximal monotone
operators. Evolution Equations \& Control Theory, (2021).


\bibitem{S} M. Schatzman, A class of nonlinear differential equations of
second order in time, Nonlinear Anal-Theor. 2 (1978).



\bibitem{SAM} F. Selamnia,  D. Azzam-Laouir,  M.D.P. Monteiro Marques,
 Evolution problems involving state-dependent maximal monotone
operators. Appl. Anal. (2020)
https://doi.org/10.1080/00036811.2020.1738401




\bibitem{Thi3} L. Thibault, Moreau sweeping process with bounded retraction. J. Convex Anal. 23,  4 (2016)
1051-1098.

%\bibitem{TOL} A.A. Tolstonogov, Compactness of BV solutions of a convex sweeping
%process of measurable differential inclusion. J. Convex Anal. 27, 2
%(2020) 673-695.

\bibitem{TOL2} A.A. Tolstonogov, BV solutions of a convex sweeping
process with a composed perturbation. Evolution Equations \& Control
theory. (2021).

\bibitem{TOL1} A.A. Tolstonogov, BV continuous solutions of an
evolution inclusion with maximal monotone operator and non-convex
valued perturbation. Existence theorem. Set-Valued Var. Anal. 29
(2021) 29-60.

%\bibitem{Va2} M. Valadier, Une propri\'et\'e de l ensemble des  s\'elections \`a variation born\'ee
%d'une multiplication a r\'etraction born\'ee. S\'em. Anal. Convexe.
%Montpellier, Expos\'e 13 (1977).


\bibitem{Vla} A.A. Vladimirov, Nonstationnary dissipative evolution equation in Hilbert space. Nonlinear Anal. 17 (1991) 499-518

\bibitem{Vra}
I.I. Vrabie, Compactness methods for nonlinear evolution equations.
Pitman Monographs and Surveys in Pure and Applied mathematics,
Longman Scientific and Technical, John Wiley  and Sons, Inc. New
York, Vol 32  (1987).

 \end{thebibliography}

\UseRawInputEncoding
\documentclass[a4paper,11pt]{article}
\usepackage{graphicx}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{amstext}
\usepackage{color}
\usepackage{colortbl}
\usepackage{xcolor}

\newcommand{\proof}{\vskip 5truemm \noindent{\textsc Proof.}~}

\newtheorem{rem}{Remark}[section]
\newtheorem{thm}{Theorem}[section]
\newtheorem{coro}{Corollary}[section]
\newtheorem{lem}{Lemma}[section]

\numberwithin{equation}{section}
\begin{document}

\title{\bf BV right-continuous solutions of second-order differential inclusions governed by maximal monotone operators}

\author{Dalila Azzam-Laouir
 \footnote{LAOTI, FSEI, Universit\'e Mohamed Seddik Benyahia de Jijel, Alg\'erie. E-mail: laouir.dalila@gmail.com,
 dalilalaouir@univ-jijel.dz}}

\maketitle

\begin{abstract} This paper concerns  existence
of right-continuous with bounded variation solutions of a perturbed
second-order differential inclusion governed by time and
state-dependent maximal monotone operators.
\end{abstract}

\vskip4mm \noindent{\footnotesize\textbf{Keywords:} Bounded
variation, differential measure, Lipschitz mapping,  maximal
monotone operator, pseudo-distance, right-continuous, second-order.}

\vskip4mm \noindent{\footnotesize\textbf{AMS Subject
Classifications: 2010}: 34A60,  28A25, 28C20}

\vskip6mm

\section{Introduction}
Let $I=[0,T]$ ($T>0$) and $\mathcal{H}$ a separable Hilbert space.
Consider, for each $(t,x)\in I\times \mathcal{H}$, a maximal
monotone operator of $\mathcal{H}$; $A(t,x)$, for which  the
set-valued map $t \mapsto A(t,x) $ is right-continuous with bounded
variation (BVRC) w.r.t the time $t$, and Lipschitz continuous w.r.t
the state variable $x$,
  in the sense  that
 there exists a function  $\rho:  I \rightarrow [0, \infty[$,
which is right-continuous on $[0, T[$,   and a nonnegative real
number $\gamma$ such that for $0\leq s \leq t \leq T,\;x, y\in
\mathcal{H}$,
$$ dis(A(t,x), A(s,y)) \leq d\rho(]s, t])+ \gamma\|x-y\|= \rho(t)-\rho(s)+\gamma\|x-y\|, $$
where  $dis(\cdot, \cdot)$ is the pseudo-distance between maximal
monotone operators introduced by Vladimirov \cite{Vla}; see relation
\eqref{2.1}, and $f:I\times \mathcal{H}\times
\mathcal{H}\longrightarrow \mathcal{H}$ a   Carath\'eodory mapping
satisfying a linear growth condition.

Let $\lambda$ be the Lebesgue measure on $I$ and $d\rho$  the
Stieljes-measure associated with $\rho$.  We set $\nu := \lambda
+d\rho$ and $\frac{ d\lambda}{d\nu}$ the density of $ \lambda$ w.r.t
 $\nu$.

 In this paper, we are mainly interested by the existence of
  BVRC  mappings $u, v: I\longrightarrow \mathcal{H}$
 satisfying
$$
(P_f)
\begin{cases}
u(0)=u_0, v(0)=v_0\in D(A(0, u_0));\\
 v(t)\in D(A(t, u(t)))\;\;\;\forall t\in I;\\
\displaystyle\frac{du}{d\nu }(t)=v(t)\;\;\;d\nu-a.e.\,t\in
   I;\\
   -\displaystyle\frac{dv}{d\nu }(t)\in A(t,u(t)) v(t)+ f(t,u(t), v(t))  \frac{d\lambda} {d\nu}(t)  \;\;\;d\nu-a.e.\,t\in
   I.
\end{cases}
$$
This work constitutes an extension to the second-order of few
results dealing with first-order BVRC time-dependent maximal
monotone operators differential inclusions, established in \cite{Az,
ACM1, KM}.

For some investigations on first-order differential inclusions
governed by time-dependent and time and state dependent maximal
monotone operators, where the variation of these operators is
absolutely continuous or continuous with bounded variation; we refer
for instance to \cite{ABCM1, ABCM2, AB, ACM, BAC, CIT, Ken, KM, Le,
P, SAM, TOL1, Vla}. Concerning the second order, we refer to the
recent works \cite{CMRF, Saidi1}.

It is well known that the normal cone to closed and convex sets is a
maximal monotone operator, so  differential inclusions governed by
this normal cone, called "sweeping processes" (\cite{M, M1, Mon})
are particular cases of the evolutions with general maximal monotone
operators.  Existence of  BVRC solutions to the first-order sweeping
process
 was   discussed in many interesting papers; we cite
for instance \cite{AHT, ANT, BCG, BCSS, CM, ET1, M2, NNT, Thi3,
TOL2}. However, to the best of our knowledge, there are no results
for the BVRC second-order sweeping process in comparison with the
vast literature concerning absolutely continuous and continuous with
bounded variation solutions; see \cite{AL, AL1, AN, AA1, AA2, AACM,
Az1, AI, Bou, BL, C1, CIY, NNT, N, S}.

The paper is organized as follows. In section 2, we give notations
and recall the needed preliminary results. In section 3, we prove
the existence   of   right-continuous with bounded variation
solutions to the evolution problem $(P_f)$ when $t \mapsto
A(t,\cdot) $ is BVRC and the perturbation $f: I\times
\mathcal{H}\times \mathcal{H}\longrightarrow \mathcal{H}$ is
measurable on $I$ and Lipshitz-continuous on  $\mathcal{H}\times
\mathcal{H}$.

\section{Notations and  Preliminaries} Throughout the paper, we will denote by $\langle \cdot ,\cdot \rangle$ the inner product
of $\mathcal{H}$,   $\|\cdot\|$ the associated norm and
$r\overline{B}_{\mathcal{H}}$ the closed
  ball of center $0$ and radius $r>0$. We will denote by $\mathcal{B}(I)$
the Borel $\sigma$-algebra on $I$. The identity mapping of
$\mathcal{H}$ will be denoted by $Id_{\mathcal{H}}$. For a subset
$K$ of $\mathcal{H}$, $\overline{co}(K)$ will be the closed convex
hull of $K$, which is characterized by:
\begin{equation}\label{co}
\overline{co}(K)=\big\{y\in \mathcal{H}:\;\langle y, x\rangle\leq
\sup_{z\in K}\langle z, x\rangle\;\forall x\in K\big\}.
\end{equation}


If $\mu$ is a positive measure on $I$, we will denote by $L^p(I,
\mathcal{H}; \mu)$ $p \in [1, +\infty[$, the Banach space of classes
of $p$-$\mu$-integrable mappings, equipped with its classical norm
$\|\cdot\|_p$.
\vskip2mm The following results concern  definitions and some
properties of functions with bounded variation and general vector
measures, taken from references \cite{Mon, M2, M3, MV}.

Let $u:I\longrightarrow \mathcal{H}$. The variation of $u$
 in $I$ is the nonnegative extended real number
 \begin{equation}\label{1}
 var(u ; I):=\sup\sum_{k=1}^{n}\|u(t_k)-u(t_{k-1})\|,
 \end{equation}
 where the supremum is taken in $[0,+\infty]$ w.r.t all the finite
 sequences $t_0<t_1<\cdots<t_n$ of points of $I$ ($n$ is arbitrary).
 The function $u$ is said with bounded variation (BV) if and only if
 $var(u ; I)<+\infty$. In this
 case we have
 \begin{equation*}
 \lim_{s\uparrow t}\|u(t)-u(s)\|=\lim_{s\uparrow t} var(u; [s,t]).
 \end{equation*}

 Now, let $\mu$ be a positive Radon measure on $I$ and $\hat{\mu}$
 be an $\mathcal{H}$-valued measure on
 $I$ admitting a density $\frac{d \hat{\mu}}{d\mu}\in
 L^1_{loc}(I,\mathcal{H};\mu)$. Then for $d \mu$-almost every $t\in
 I$, we have
 \begin{equation}\label{2}
\frac{d \hat{\mu}}{d\mu}(t)=\lim_{\varepsilon\downarrow 0}\frac{d
\hat{\mu}([t, t+\varepsilon])}{d\mu([t,
t+\varepsilon])}=\lim_{\varepsilon\downarrow 0}\frac{d
\hat{\mu}([t-\varepsilon, t])}{d\mu([t-\varepsilon, t])}.
 \end{equation}
 The measure $\hat{\mu}$ is absolutely continuous w.r.t $\mu$ if and
 only if $\hat{\mu}=\frac{d \hat{\mu}}{d\mu}\mu$, i.e.,  $\frac{d
 \hat{\mu}}{d\mu}$ is a density of $\hat{\mu}$  w.r.t $\mu$. In this
 case a mapping $u:I\longrightarrow \mathcal{H}$ is
 $\hat{\mu}$-integrable if and only if the mapping $t\mapsto u(t)\frac{d
 \hat{\mu}}{d\mu}(t)$ is $\mu$-integrable, and we have
 \begin{equation}\label{7}
 \int_I u(t) d\hat{\mu}(t)=\int_I u(t)\frac{d
 \hat{\mu}}{d\mu}(t) d\mu(t).
 \end{equation}
If $u:I\longrightarrow \mathcal{H}$ is BVRC and $du$ is its
 differential measure, then we have
 \begin{equation}\label{3}
 u(t)=u(s)+\int_{]s,t]} du\;\;\;\;\;\forall s, t\in I\;(s\leq t).
 \end{equation}
 Conversely, if there exists $v\in L^1(I,\mathcal{H};\mu)$ such that
 $u(t)=u(0)+\int_{]0, t]} v \,d \mu$ for all $t\in I$, then $u$ is
 BVRC and $du=v \, d\mu$, that is $v$ is a density of the vector
 measure $du$ w.r.t the measure $\mu$. So that by \eqref{2}, we get
 \begin{equation}\label{4}
v(t)=\frac{d u}{d\mu}(t)=\lim_{\varepsilon\downarrow 0}\frac{ du
([t, t+\varepsilon])}{d\mu([t,
t+\varepsilon])}=\lim_{\varepsilon\downarrow 0}\frac{ du
([t-\varepsilon, t])}{d\mu([t-\varepsilon, t])}.
 \end{equation}
 If $\mu(\{t\})>0$, this last relation shows that
 \begin{equation}\label{5}
 v(t)=\frac{d u}{d\mu}(t)=\frac{d
 u(\{t\})}{d\mu(\{t\})}\;\;\;\textmd{and}\;\;\;\frac{d
 \lambda}{d\mu}(t)=0.
 \end{equation}

Next, we give definition and some properties of maximal monotone
operators. We refer the reader to \cite{Ba, brezis, Vra} for more
details.

 Let $A: \mathcal{H}
\rightrightarrows \mathcal{H}$ be a set-valued map. We denote by
$D(A)$ and $Gr(A)$ its domain  and graph. We say that the operator
$A$ is monotone, if $\langle y_1 -y_2, x_1 - x_2\rangle\ge 0$ for
all $(x_i, y_i)\in Gr(A)$ $(i=1, 2)$, and
  we say that $A$ is a maximal monotone operator of $\mathcal{H}$,
  if it is monotone and
its graph could not be contained strictly in the graph of any other
monotone operator.

If $A$ is a maximal monotone operator of $\mathcal{H}$, then  for
every $x \in D(A)$, $A(x)$ is nonempty closed and convex. We denote
the projection of the origin on the set $A(x)$ by $A^0(x)$.

 For $\eta > 0$, we denote by $J_{\eta}^A= (Id_{\mathcal{H}} +
\eta A)^{-1}$  the resolvent of  $A$. 
It is
 well-known that
  this operator
is single-valued and defined on all of $\mathcal{H}$, furthermore
$J^A_{\eta}(x)\in D(A)$, for all $x\in \mathcal{H}$.

Let $A: D(A)\subset \mathcal{H}\rightrightarrows \mathcal{H}$ and
$B: D(B)\subset \mathcal{H}\rightrightarrows \mathcal{H}$ be two
maximal monotone operators. Then,  $dis(A, B)$ is the Vladimirov's
pseudo-distance between $A$ and $B$, defined as follows
\begin{equation}\label{2.1} dis(A,B)=\sup\bigg\{\frac{\langle y-y', x'-x\rangle}{1+\|
y\|+\| y'\|}:\;(x,y)\in Gr(A),\;(x',y')\in
Gr(B)\bigg\}.\end{equation}

The following lemmas are  needed for our proof. We refer to
\cite{KM} for their proofs.
\begin{lem}\label{lem2.1}
Let $A$ be a maximal monotone operator of $\mathcal{H}$. If $x\in
\overline{D(A)}$ and  $y\in \mathcal{H}$ are such that
$$ \langle A^0(z)-y, z-x\rangle\geq 0\;\;\forall z\in D(A),$$
then $x\in D(A)$ and $y\in A(x)$.
\end{lem}
\begin{lem}\label{lem2.2}
Let $A_n$ $(n\in \mathbb{N})$ and $A$ be maximal monotone operators
of $\mathcal{H}$ such that $dis(A_n, A)\to 0$. Suppose also that
$x_n\in D(A_n)$ with $x_n\to x$ and $y_n\in A_n(x_n)$ with $y_n\to
y$ weakly for some $x, y\in E$. Then $x\in D(A)$ and $y\in A(x)$.
\end{lem}
\begin{lem}\label{lem2.3}
Let $A$ and $B$ be maximal monotone operators of $\mathcal{H}$. Then\\
1) for $\eta>0$ and $x\in D(A)$
$$ \|x-J_{\eta}^{B}(x)\| \leq \eta\|A^0(x)\|+dis(A,B)+\sqrt{\eta\big(1+\|A^0(x)\|\big)dis(A, B)}.$$
2) For $\eta>0$ and $x, x'\in E$
$$ \|J_{\eta}^{A}(x)-J_{\eta}^{A}(x')\|\leq
\|x-x'\|.$$
\end{lem}
\begin{lem}\label{lem2.4}
Let $A_n$ $(n\in \mathbb{N})$ and $A$ be maximal monotone operators
of $\mathcal{H}$ such that $dis(A_n, A)\to 0$ and $\|A^0_n(x)\|\leq
c(1+\|x\|)$ for some $c>0$, all $n\in \mathbb{N}$ and $x\in D(A_n)$.
Then for every $z\in D(A)$ there exists a sequence $(\zeta_n)$ such
that
 \begin{equation*}\label{2.2}\zeta_n\in
D(A_n),\;\;\;\zeta_n\to z\;\;\textmd{and}\;\;A_n^0(\zeta_n)\to
A^0(z).\end{equation*}
\end{lem}

We end this section by the following discrete Gronwall's lemma.

\begin{lem}\label{lem2.5}
Let $(\alpha_i)$, $(\beta_i)$, $(\gamma_i)$ and $(a_i)$ be sequences
of nonnegative real numbers such that $a_{i+1}\leq
\alpha_i+\beta_i\big(a_0+a_1+\cdots+a_{i-1}\big)+(1+\gamma_i)a_i$
for $i\in \mathbb{N}$. Then
$$ a_j\leq \bigg(a_0+\sum_{k=0}^{j-1}
\alpha_k\bigg)\exp\bigg(\sum_{k=0}^{j-1}
\big(k\beta_k+\gamma_k\big)\bigg)\;\;\textmd{for}\;j\in
\mathbb{N}^*.$$
\end{lem}
\vskip2mm

We will establish our main result under the following hypotheses.\\
$(H_1)$ There exist a function $\rho:I\longrightarrow [0, +\infty[$,
which is right-continuous on $[0, T[$ and nondecreasing with
$\rho(0) = 0$ and $\rho(T)<+\infty$, and nonnegative constant
$\gamma$ such that
\begin{equation*}dis(A(t,x), A(s,y))\leq d\rho(]s,
t])+\gamma\|x-y\|\frac{d\lambda}{d\nu}(s)\;\;\textmd{for}\;\;0\leq
s\leq t\leq T,\;x,y\in \mathcal{H},\end{equation*}  and since
$\frac{d\lambda}{d\nu}(s)\leq 1$, $(H_1)$ implies that
\begin{equation*}dis(A(t,x), A(s,y))\leq d\rho(]s,
t])+\gamma\|x-y\|\;\;\textmd{for}\;\;0\leq s\leq t\leq T,\;x,y\in
\mathcal{H}.\end{equation*}
 $(H_2)$ There
exists a nonnegative real constant $c$ such that
\begin{equation*}\|A^0(t,x)y\|\leq c(1+\| x\|+\| y\|)\;\;\textmd{for}\;\;t\in
I,\;x\in\mathcal{H},\;y\in D(A(t,x)).\end{equation*}
 $(H_3)$ For any bounded subset $E$ of $\mathcal{H}$, $D(A(I\times E))$
 is relatively ball-compact, i.e., its intersection with any closed ball of
 $\mathcal{H}$ is relatively compact.\\
$(H_4)$ $t\mapsto f(t,\cdot,\cdot)$ is $\mathcal{B}(I)$-measurable.\\
$(H_5)$ There  is a nonnegative function $k\in L^1(I,\mathbb{R};
\lambda)$ such that
$$ \|f(t,x,y) - f(t,x',y')\|\leq k(t)(\|x-x'\|+\|y-y'\|)\;\;\;\forall t\in I, \;\forall (x,y), (x',y')\in
\mathcal{H}\times \mathcal{H}.$$
 $(H_6)$ There exists a nonnegative real constant $m$
such that
$$ \|f(t,x,y)\|\leq m(1+\|x\|
+\|y\|)\;\;\;\forall (t,x,y)\in I\times \mathcal{H}\times
\mathcal{H}.$$

\section{Main result}
Now we are able to state our main theorem.
\begin{thm}\label{Theorem 3.1}    Let for
every $(t,x)\in I\times \mathcal{H}$, $A(t,x):D(A(t,x))\subset
\mathcal{H}\rightrightarrows \mathcal{H}$ be a maximal monotone
operator satisfying $(H_1)$, $(H_2)$ and $(H_3)$. Let $f:I\times
\mathcal{H}\times \mathcal{H}\longrightarrow \mathcal{H}$  such that
$(H_4)$, $(H_5)$ and $(H_6)$ are satisfied. Then for any
$(u_0,v_0)\in \mathcal{H}\times D(A(0, u_0))$, there exists a BVRC
solution $ (u, v) :I \longrightarrow \mathcal{H}\times \mathcal{H}$
to  problem $(P_f)$.
\end{thm}

\proof
 Following
Castaing et al \cite{CM}, we can choose a sequence
$(\varepsilon_n)_{n\geq 1}\subset ]0, 1]$ such that $\varepsilon_n
\downarrow 0$ and a partition $0=t_0^n<t_1^n<\cdots<t_{q_n}^n=T$ of
$I$ (with $\underset{n\to\infty}{\lim}q_n=+\infty$), for which we
have
\begin{equation}\label{3.1}
\nu(]t^n_i, t^n _{i+1}])=|t_{i+1}^n-t_i^n|+d\rho(]t_i^n,
t_{i+1}^n])\leq
\varepsilon_n\;\;\;\textmd{for}\;i=0,\cdots,q_n-1,\end{equation} we
can take, for example, $\varepsilon_n=\frac{T}{q_n}$. For each $i\in
\{0,\cdots,q_n-1\}$, put $I_i^n=]t_i^n, t_{i+1}^n]$,
\begin{equation}\label{3.2} \delta_{i+1}^n=d\rho(]t_i^n,
t_{i+1}^n]), \;\;\;\; \eta^n_{i+1} = t ^n_{i+1}-t ^n_i, \;\;\;\;
\beta ^n_{i+1}  = \nu(]t^n_i, t^n _{i+1}]).\end{equation} For every
$n\in \mathbb{N}$, put $u_0^n=u_0$, $v_0^n=v_0\in D(A(0,u_0))$, and
let us define by induction,  sequences $(v_i^n)_{0\leq i\leq
q_n-1}$, $(u_i^n)_{0\leq i\leq q_n-1}$ such that
\begin{equation}\label{3.3} v_{i+1}^n= J^n_{i+1}
\Big(v_i^n-\int_{t_i^n}^{t_{i+1}^n} f(s, u_i^n,
v_i^n)d\lambda(s)\Big)\end{equation}  where
$J^n_{i+1}:=J^{A(t_{i+1}^n,u_i^n )}_{\beta^n_{i+1}}
=\big(Id_{\mathcal{H}}+ \beta^n_{i+1} A(t_{i+1}^n,
u_i^n)\big)^{-1}$, and
\begin{equation}\label{3.2'}
u_{i+1}^n=u_i^n+\beta^n_{i+1}v_{i+1}^n.
\end{equation}
  From the properties of the resolvent,  we have that  $v^n_{i+1}
\in D(A(t_{i+1}^n, u_i^n))$ and
\begin{equation}\label{3.5} -\frac{1} {\beta^n_{i+1}}\Big(v^n_{i+1}-
v_i^n+\int_{t_i^n}^{t_{i+1}^n} f(s, u_i^n, v_i^n)d\lambda(s)\Big)
\in A( t_{i+1}^n, u_i^n) v^n _{i+1}. \end{equation} For $t\in
[t^n_i, t ^n_{i+1}[$,  $i = 0,\cdots,q_n-1$, set
\begin{equation}\label{3.7}
v_n(t)=   v_i^n + \frac{ \nu(]t^n_i, t])  }{ \nu(]t^n_i, t
^n_{i+1}])} \Big(v_{i+1}^n-v_i^n+\int_{t_i^n}^{t_{i+1}^n} f(s,u_i^n,
v_i^n) d\lambda(s)\Big)  - \int_{t_i^n}^{t}
f(s,u_i^n,v_i^n)d\lambda(s),\end{equation}
\begin{equation}\label{3.7'}
u_n(t)=   u_i^n +  \frac{ \nu(]t^n_i, t])  }{ \nu(]t^n_i, t
^n_{i+1}])}  (u_{i+1}^n-u_i^n),
\end{equation} and
$v_n(T)=v_{q_n}^n$, $u_n(T)=u_{q_n}^n$. So that $v_n(t_i^n)=v_i^n$,
$u_n(t_i^n)=u_i^n$ and $v_n$, $u_n$ are BVRC mappings on $I$.

\vskip2mm{\bf Step 1.}
 We prove in this step  that the sequences $(v_n)$, $(u_n)$ are bounded in  norm and
 in variation.\\Using Lemma \ref{lem2.3}, we have from
\eqref{3.3}, $(H_1)$,  $(H_2)$ and $(H_6)$, for $i=0,1,\cdots,
q_n-1$ (keeping in mind that $\sqrt{a\,b}\leq \frac{a}{2}+b$ for
nonnegative real numbers $a$, $b$)
\begin{eqnarray*}
\| v_{i+1}^n-v_i^n\| &\leq& \Big\|J_{i+1}^n\Big(v_i^n-
\int_{t_i^n}^{t_{i+1}^n}
f(s,u_i^n,v_i^n)d\lambda(s)\Big)-J_{i+1}^n(v_i^n)\Big\|+\big\|J_{i+1}^n(v_i^n)-v_i^n\big\|\\
&\leq& \int_{t_i^n}^{t_{i+1}^n} \|f(s,u_i^n,v_i^n)\|
d\lambda(s)+\beta_{i+1}^n\|A^0(t_i^n,
u_{i-1}^n)v_i^n\|+dis\big(A(t_{i+1}^n, u_i^n),A(t_i^n,
u_{i-1}^n)\big)\\&+&\sqrt{\beta_{i+1}^n\big(1+\|A^0(t_i^n,
u_{i-1}^n)v_i^n\|\big)dis\big(A(t_{i+1}^n, u_{i}^n),A(t_i^n, u_{i-1}^n)\big)}\\
&\leq& m(1+\|u_i^n\|+\|v_i^n\|)\beta_{i+1}^n
+\big(1+c(1+\|u_{i-1}^n\|+\|v_i^n\|)\big)\beta_{i+1}^n+\gamma\|u_{i}^n-u_{i-1}^n\|\\&+&
\sqrt{\beta_{i+1}^n\big(1+c(1+\|u_{i-1}^n\|+\|v_i^n\|)\big)(\beta_{i+1}^n+\gamma\|u_{i}^n-u_{i-1}^n\|)}\\
&\leq& m(1+\|u_i^n\|+\|v_i^n\|)\beta_{i+1}^n
+\big(1+c(1+\|u_{i-1}^n\|+\|v_i^n\|)\big)\beta_{i+1}^n+\gamma\|u_{i}^n-u_{i-1}^n\|\\
&+&\beta_{i+1}^n\frac{1}{2}\big(1+c(1+\|u_{i-1}^n\|+\|v_i^n\|)\big)+\beta_{i+1}^n+\gamma\|u_{i}^n-u_{i-1}^n\|,
\end{eqnarray*}
that is, \begin{eqnarray}\label{3.8}\| v_{i+1}^n-v_i^n\|&\leq&
(m+1)(1+\|u_i^n\|+\|v_i^n\|)\beta_{i+1}^n+\frac{3}{2}(1+c)(1+\|u_{i-1}^n\|+\|v_{i}^n\|)\beta_{i+1}^n+2\gamma\|u_{i}^n-u_{i-1}^n\|\nonumber\\&=:&
m_1(1+\|u_i^n\|+\|v_i^n\|)\beta_{i+1}^n+c_1(1+\|u_{i-1}^n\|+\|v_{i}^n\|)\beta_{i+1}^n+\gamma_1\|u_{i}^n-u_{i-1}^n\|,\end{eqnarray}
which entails \begin{equation} \label{eq0}\|v_{i+1}^n\|\leq
\|v_i^n\|+m_1(1+\|u_i^n\|+\|v_i^n\|)\beta_{i+1}^n+c_1(1+\|u_{i-1}^n\|+\|v_{i}^n\|)\beta_{i+1}^n+\gamma_1\|u_{i}^n-u_{i-1}^n\|
.\end{equation} On the other hand, we have from \eqref{3.2'},
\begin{eqnarray}\label{eq1}
\|u_{i}^n-u_{i-1}^n\|=\beta_i^n\|v_i^n\|
\end{eqnarray}
and (using \eqref{3.1})
\begin{eqnarray}\label{eq2}
\|u_{i}^n\|&=&\|u_{i-1}^n+\beta_i^nv_i^n\|=\|u_0+\beta_1^nv_1^n+\beta_2^nv_2^n+\cdots+\beta_i^nv_i^n\|\nonumber\\&\leq&
\|u_0\|+\varepsilon_n(\|v_1^n\|+\|v_2^n\|+\cdots+\|v_i^n\|).
\end{eqnarray}
Replacing  these two last relations in \eqref{eq0}, we get by the
use of \eqref{3.1},
\begin{eqnarray*}
\|v_{i+1}^n\|&\leq&\|v_i^n\|+m_1\Big(1+\|u_0\|+\varepsilon_n(\|v_1^n\|+\|v_2^n\|+\cdots+\|v_i^n\|)+\|v_i^n\|\Big)\varepsilon_n\\&+&
c_1\big(1+\|u_0\|+\varepsilon_n(\|v_1^n\|+\|v_2^n\|+\cdots+\|v_{i-1}^n\|)+\|v_i^n\|\big)\varepsilon_n+\gamma_1\|v_i^n\|\varepsilon_n\\&\leq&
(m_1+c_1)(1+\|u_0\|)\varepsilon_n+(m_1+c_1)(\|v_1^n\|+\|v_2^n\|+\cdots+\|v_{i-1}^n\|)\varepsilon^2_n
\\&+&\Big(1+(m_1\varepsilon_n+m_1+c_1+\gamma_1)\varepsilon_n\Big)\|v_i^n\|.
\end{eqnarray*}
Applying Lemma \ref{lem2.5}, we get for all $i\leq q_n$,
\begin{eqnarray*}\|v_i^n\|&\leq&
\Big(\|v_0\|+\sum_{j=0}^{i-1}
(m_1+c_1)(1+\|u_0\|)\varepsilon_n\Big)\exp\Big(\sum_{j=0}^{i-1}
\big(j(m_1+c_1)\varepsilon^2_n+(m_1\varepsilon_n+m_1+c_1+\gamma_1)\varepsilon_n\big)\Big)\\
&\leq&\Big(\|v_0\|+(m_1+c_1)(1+\|u_0\|)T)\Big)\exp\Big((2+T)m_1+2c_1+\gamma_1
\Big)=:m_2,
\end{eqnarray*}
using this last estimate in \eqref{eq2}, we obtain
$$\|u_i^n\|\leq \|u_0\|+m_2 T,$$ and replacing in
 \eqref{eq1} and in \eqref{3.8}, it comes that
 $$\|u_{i+1}^n-u_i^n\|\leq m_2\beta_{i+1}^n,$$ and
$$\|v_{i+1}^n-v_i^n\|\leq(m_1+c_1) (1+\|u_0\|+m_2T+m_2)\beta_{i+1}^n+\gamma_1m_2\beta_{i}^n=:m_3\beta_{i+1}^n+m_4\beta_{i}^n.$$
Consequently,
\begin{equation}\label{3.9}\|u_i^n\|\leq M,\;\;\;\|v_i^n\|\leq M\;\;\forall\, 0\leq i \leq
q_n\end{equation} and for all $0\leq i<q_n-1$,
 \begin{equation}\label{3.9'}\|u_{i+1}^n-u_i^n\|\leq M\nu(]t_i^n,
t_{i+1}^n]),\;\;\;\|v_{i+1}^n-v_i^n\|\leq M\big(\nu(]t_i^n,
t_{i+1}^n])+\nu(]t_{i-1}^n, t_{i}^n])\big),
\end{equation}
where $M=\max(m_2, m_3, m_4)$. Now, for $t\in [t_i^n, t_{i+1}^n[$,
we have from \eqref{3.1}, \eqref{3.7}, \eqref{3.9}, \eqref{3.9'} and
$(H_6)$,
\begin{equation*}\label{3.10}
\|v_n(t)-v_i^n\| \leq
2M\varepsilon_n+2m(1+2M)\varepsilon_n\leq2M+2m(1+2M)=:M_1,\end{equation*}
and from \eqref{3.1}, \eqref{3.7'},  \eqref{3.9} and \eqref{3.9'},
we get
\begin{equation*}\label{3.10'}
\|u_n(t)-u_i^n\| \leq M.
\end{equation*}
These last relations give us, for all $n\in \mathbb{N}$,
\begin{equation*}\|v_n(t)\|\leq M+M_1=:M_2\;\;\;\textmd{and}\;\;\;\|u_n(t)\|\leq 2M\;\;\;\forall t\in I,\end{equation*}  that is
\begin{equation*}\label{3.11}
\sup_n\|v_n\|:=\sup_n\big(\sup_{t\in I}\|v_n(t)\|\big)\leq
M_2\;\;\;\textmd{and}\;\;\;\sup_n\|u_n\|\leq 2M.
\end{equation*}
That is $(u_n)$, $(v_n)$ are bounded in norm and in variation.
\vskip2mm

{\bf Step 2.} Convergence of the sequences $(u_n)$, $(v_n)$ and $(\frac{dv_n}{d\nu})$.\\
Define the functions $\phi_n,\;\theta_n: I\longrightarrow I$ by
$$\phi_n(t)=t_{i}^n,\;\;\;\theta_n(t)=t_{i+1}^n\;\;\textmd{for}\;\; t\in ]t_i^n,
t_{i+1}^n],\;i=0,1,\cdots,q_n-1,\;\; \textmd{and}\;\;
\phi_n(0)=\theta_n(0)=0,
$$ and set for all $t\in I$,
\begin{equation*}
B_n(t)=\sum_{i=0}^{q_n-1}\frac{1}{\nu(]t_i^n, t_{i+1}^n])}\Big(
v_{i+1}^n-v_i^n+\int_{t_i^n}^{t_{i+1}^n} f(s,u_i^n,
v_i^n)d\lambda(s)\Big)\chi_{]t_i^n, t_{i+1}^n]}(t),
\end{equation*}
where $\chi_{J}(\cdot)$ is the characteristic function of the set
$J\subset I$. Whence, by \eqref{3.7}, it is clear that
\begin{equation*}
v_n(t)=v_0+\int_{]0,t]} B_n(\tau)d\nu(\tau)-\int_{]0,t]} f(\tau,
u_n(\phi_n(\tau)), v_n(\phi_n(\tau)))d\lambda(\tau).
\end{equation*}
Since $\frac{d\lambda}{d\nu}$ is a density of $\lambda$ w.r.t $\nu$,
 then by \eqref{7}, we get for every $t\in I$,
\begin{equation*}
v_n(t)=v_0+\int_{]0,t]} \Big(B_n(\tau)-f(\tau, u_n(\phi_n(\tau)),
v_n(\phi_n(\tau)))\frac{d\lambda}{d\nu}(\tau)\Big)d\nu(\tau),
\end{equation*}
so that, $\frac{dv_n}{d\nu}$ is a density of the vector measure
$dv_n$ w.r.t $\nu$ and for $\nu$-almost every $t\in I$,
\begin{equation*}\label{8}\frac{dv_n}{d\nu}(t)=B_n(t)-f(t,
u_n(\phi_n(t)),
v_n(\phi_n(t)))\frac{d\lambda}{d\nu}(t),\end{equation*} that is, by
the definition of $B_n$, for $\nu$-almost every $t\in I$,
\begin{eqnarray}\label{8}
&&\frac{dv_n}{d\nu}(t)+f(t,
u_n(\phi_n(t)),v_n(\phi_n(t)))\frac{d\lambda}{d\nu}(t)\nonumber\\&=&\sum_{i=0}^{q_n-1}\frac{1}{\nu(]t_i^n,
t_{i+1}^n])}\Big( v_{i+1}^n-v_i^n+\int_{t_i^n}^{t_{i+1}^n}
f(s,u_i^n, v_i^n)d\lambda(s)\Big)\chi_{]t_i^n, t_{i+1}^n]}(t).
\end{eqnarray}
 From $(H_6)$, \eqref{3.9} and \eqref{8}, it results that
 \begin{equation}\label{9}
 \Big\|\frac{dv_n}{d\nu}(t)+f(t,
u_n(\phi_n(t)), v_n(\phi_n(t)))\frac{d\lambda}{d\nu}(t)\Big\|\leq
M+m(1+2M)=:M_3\;\;\;\nu-a.e.\,t\in I.
 \end{equation}
On the other hand, we know that $\nu=d\rho+\lambda$, whence
$\frac{d\lambda}{d\nu}(t)\leq 1$. So that, by $(H_6)$ and
\eqref{3.9}, it follows that
\begin{equation}\label{10}
 \Big\|f(t,
u_n(\phi_n(t)), v_n(\phi_n(t)))\frac{d\lambda}{d\nu}(t)\Big\|\leq
m(1+2M)=:M_4\;\;\;\nu-a.e.\,t\in I,
 \end{equation}
 consequently, there is a Borel subset $J\subset I$ with $\nu(J)=0$,
 such that
 \begin{equation}\label{11}
 \Big\|\frac{dv_n}{d\nu}(t)\Big\|\leq M_3+M_4=:M_5\;\;\;\forall t\in I\setminus J.
 \end{equation}
 Now, observe that by \eqref{3.5} and \eqref{8}, for each $n\in \mathbb{N}$, there is a
 Borel subset $J_n\subset I$ with $\nu(J_n)=0$ such that
 \begin{equation}\label{12}
 \frac{dv_n}{d\nu}(t)+f(t,
u_n(\phi_n(t)),v_n(\phi_n(t)))\frac{d\lambda}{d\nu}(t)\in
-A(\theta_n(t), u_n(\phi_n(t)))v_n(\theta_n(t))\;\;\;\forall t\in
I\setminus J_n,
\end{equation}
moreover,
\begin{equation}\label{13}
v_n(\theta_n(t))\in
D\big(A(\theta_n(t),u_n(\phi_n(t)))\big)\;\;\;\forall t\in I.
\end{equation}

We will show in the following, that $(v_n)$ converges uniformly on $I$.\\
On the one hand, since $\frac{dv_n}{d\nu}$ is a density of the
vector measure $dv_n$ w.r.t the measure $\nu$, we have from
\eqref{3}, for all $s, t\in I$ $(s\leq t)$,
\begin{eqnarray*}
v_n(t)=v_n(s)+\int_{]s, t]} \frac{dv_n}{d\nu}(s)\,d\nu(s),
\end{eqnarray*}
that is, using relation \eqref{11}, we get
\begin{equation}\label{eq3}\|v_n(t)-v_n(s)\|\leq M_5 \nu(]s,
t]),\end{equation} which means that $(v_n)$ is equi-right-continuous
with bounded variation. On the other hand, from \eqref{3.9} and
\eqref{13}, it is clear that for any $t\in I$, we have
$$\big(v_n(\theta_n(t))\big)\subset D(I\times
M\overline{B}_{\mathcal{H}})\cap M \overline{B}_{\mathcal{H}},$$
which shows that the sequence $\big(v_n(\theta_n(t))\big)$ is
relatively compact (see $(H_3))$. From \eqref{eq3}, we have that for
all $t\in I$,
\begin{equation}\label{eq4}
\|v_n(\theta_n(t))-v_n(t)\|\leq M_5d\nu(]t, \theta_n(t)])\to
0\;\;\textmd{as}\;\;n\to\infty,
\end{equation}
whence, we conclude that $(v_n(t))$ is also relatively compact.
Using Helly-Banach's Theorem (see \cite{Por}), we may assume the
existence of a BVRC mapping $v$ such that $(v_n(t))$ converges
strongly to $v(t)$ for every $t\in I$. This together with
\eqref{eq3}, show that
\begin{equation}\label{eq3'}\|v(t)-v(s)\|\leq M_5 \nu(]s,
t]),\end{equation} and from \eqref{eq4}
\begin{equation}\label{eq4'}
\|v_n(\theta_n(t))-v(t)\|\leq
\|v_n(\theta_n(t))-v_n(t)\|+\|v_n(t)-v(t)\|\to
0\;\;\textmd{as}\;\;n\to\infty.
\end{equation}
As regards the sequence $(u_n)$, let us set for all $t\in I$,
\begin{equation*}
\alpha_n(t)=\sum_{i=0}^{q_n-1}\frac{1}{\nu(]t_i^n, t_{i+1}^n])}
(u_{i+1}^n-u_i^n)\chi_{]t_i^n, t_{i+1}^n]}(t),
\end{equation*}
and remark that by relation \eqref{3.2'},
$\alpha_n(t)=v_n(\theta_n(t))$ for $\nu$-a.e.  $t\in I$.
Consequently, from \eqref{3.7'}, it comes that for all $t\in I$,
\begin{equation}\label{eq5}
u_n(t)=u_0+\int_{]0, t]}\alpha_n(\tau) d\nu(\tau)=u_0+\int_{]0,
t]}v_n(\theta_n(\tau))d\nu(\tau).
\end{equation}
Using relations \eqref{3.9} and \eqref{eq4} and the strong
convergence in $\mathcal{H}$ of $(v_n(t))$ to $v(t)$, we obtain by
Lebesgue dominated convergence, that
$$ \lim_{n\to\infty} u_n(t)=u_0+\int_{]0,
t]}v(\tau)d\nu(\tau)=:u(t)$$ so that, $u$ is a BVRC mapping,
$(u_n(t))$ converges strongly to  $u(t)$ for all $t\in I$, and
$\frac{du}{d\nu}$ is a density of the vector measure $du$ w.r.t
$\nu$ and for $\nu$-almost every $t\in I$,
\begin{equation}\label{noter}
\frac{du}{d\nu}(t)=v(t).
\end{equation}

Now, observe that the sequence $(\frac{dv_n}{d\nu})$ is bounded in
$L^2(I, \mathcal{H}; \nu)$ due to relation \eqref{11}, so that it
converges weakly in $L^2(I, \mathcal{H}; \nu)$ to some mapping $w\in
L^2(I, \mathcal{H}; \nu)$. In particular, for all $t\in I$
\begin{equation*}
\int_{]0, t]} \frac{dv_n}{d\nu}(s)\, d\nu(s)\longrightarrow\int_{]0,
t]} w(s)\, d\nu(s)\;\;\;\textmd{weakly in }\;\mathcal{H}.
\end{equation*}
Since $\frac{dv_n}{d\nu}$ is a density of the vector measure $dv_n$
w.r.t the measure $\nu$, we have for all $t\in I$,
\begin{eqnarray*}
v_n(t)=v_0+\int_{]0, t]} \frac{dv_n}{d\nu}(s)\,d\nu(s)
\end{eqnarray*}
and since for any $t\in I$, $(v_n(t))$ converges strongly in
$\mathcal{H}$ and then weakly to $v(t)$, we deduce by what precedes
that
\begin{equation*}
v(t)=v_0+\int_{]0, t]} w(s)\,d\nu(s)\;\;\;\forall t\in I.
\end{equation*}
This shows that $\frac{dv}{d\nu}=w$ for $\nu$-a.e. $t\in I$.
Consequently, $(\frac{dv_n}{d\nu})$ converges weakly in $L^2(I,
\mathcal{H}; \nu)$ to $\frac{dv}{d\nu}$.

\vskip2mm

{\bf Step 3.} Existence of solution.\\
First, observe that from \eqref{13},  $ v_n(\theta_n(t))\in
D\big(A(\theta_n(t),u_n(\phi_n(t)))\big)$ for all $t\in I $. Also,
from $(H_1)$, we have that for all $t\in I$,
$$dis\big(A(\theta_n(t), u_n(\phi_n(t))), A(t, u(t))\big)\leq
|\rho(\theta_n(t))-\rho(t)|+\gamma\|u_n(\phi_n(t))-u(t)\|\frac{d\lambda}{d\nu}(t),$$
but from \eqref{eq5}, it is clear that
$$\|u_n(\phi_n(t))-u(t)\|\leq \|u_n(\phi_n(t))-u_n(t)\|+\|u_n(t)-u(t)\|\leq M\nu(]\phi_n(t), t])+\|u_n(t)-u(t)\|,$$
so, from \eqref{5}, it comes that
\begin{equation}\label{eq5'}\|u_n(\phi_n(t))-u(t)\|\frac{d\lambda}{d\nu}(t)\leq
\big(M\nu(]\phi_n(t),
t])+\|u_n(t)-u(t)\|\big)\frac{d\lambda}{d\nu}(t)\longrightarrow
0\;\textmd{as}\;n\to\infty.\end{equation} This shows that
\begin{equation}\label{eq6}
dis\big(A(\theta_n(t), u_n(\phi_n(t))), A(t,
u(t))\big)\longrightarrow 0\;\textmd{as}\;n\to\infty.
\end{equation}
On the other hand, by \eqref{3.9} and $(H_2)$, we have that the
sequence $\big(A^0\big(\theta_n(t),
u_n(\phi_n(t))\big)v_n(\theta_n(t))\big)$ is bounded in
$\mathcal{H}$, and hence it is weakly relatively compact. Using this
fact, \eqref{eq4'} and \eqref{eq5'}, we conclude by Lemma
\ref{lem2.2}, that $v(t)\in D(A(t, u(t)))$ for all $t\in I$.

Next, remember that the sequence $(\frac{dv_n}{d\nu})$ converges
weakly in $L^2(I, \mathcal{H}; \nu)$ to $\frac{dv}{d\nu}$, so that
by Mazur's theorem, there is a sequence $(\xi_n)$ such that for each
$n\in\mathbb{N}$, $\xi_n\in co\{\frac{dv_k}{d\nu};\;k\geq n\}$, and
$(\xi_n)$ converges strongly in $L^2(I, \mathcal{H}; \nu)$ to
$\frac{dv}{d\nu}$. Whence, there is a subsequence $(\xi_{n_j})$,
which converges $\nu$-almost every where to $\frac{dv}{d\nu}$. This
means the existence of a Borel subset $J'\subset I$, with
$\nu(J')=0$ and for $t\in I\setminus J'$,
$$\xi_{n_j}(t)\longrightarrow \frac{dv}{d\nu}(t)\in\bigcap_{n}
\overline{co} \big\{\frac{dv_{k}}{d\nu};\;k\geq n\big\}.$$ This
implies, by \eqref{co}, that for any fixed $\eta\in \mathcal{H}$,
\begin{equation}\label{17}
\big\langle \frac{dv}{d\nu}(t), \eta
\big\rangle\leq\limsup_{n\to\infty}\big\langle \frac{dv_n}{d\nu}(t),
\eta \big\rangle.
\end{equation}
On the other hand, we have from $(H_5)$ and \eqref{eq3}
\begin{eqnarray*}
&&\big\|f\big(t,u_n(\phi_n(t)),v_n(\phi_n(t))\big)\frac{d\lambda}{d\nu}(t)-f(t,u(t),v(t))\frac{d\lambda}{d\nu}(t)\big\|
\\&\leq&k(t)\big(\|u_n(\phi_n(t))-u(t)\|+\|v_n(\phi_n(t))-v(t)\|\big)\frac{d\lambda}{d\nu}(t)\nonumber\\
&\leq&k(t)\big(\|u_n(\phi_n(t))-u(t)\|+\|v_n(\phi_n(t))-v_n(t)\|+\|v_n(t)-v(t)\|\big)\frac{d\lambda}{d\nu}(t)\nonumber\\
&\leq&k(t)\big(\|u_n(\phi_n(t))-u(t)\|+M_5\nu(]\phi_n(t),
t])+\|v_n(t)-v(t)\|\big)\frac{d\lambda}{d\nu}(t),
\end{eqnarray*}
whence by the  convergence of $(v_n(t))$ to $v(t)$ and relations
 \eqref{5}, \eqref{eq5'}, it results that for all $t\in I$,
\begin{equation}\label{lip}
f\big(t,u_n(\phi_n(t)),
v_n(\phi_n(t))\big)\frac{d\lambda}{d\nu}(t)\longrightarrow f(t,u(t),
v(t))\frac{d\lambda}{d\nu}(t)\;\;\;\textmd{as}\;\;n\to\infty.
\end{equation}

Now, to prove that $(u, v)$ is a solution to our considered problem,
we will use Lemma \ref{lem2.1}. Since $v(t)\in D(A(t,u(t)))$ for all
$t\in I$, we have to show that for $\nu$-almost every fixed $t\in I$
and for any $z\in D(A(t, u(t)))$
\begin{equation*}
\big\langle
A^0(t,u(t))z+\frac{dv}{d\nu}(t)+f(t,u(t),v(t))\frac{d\lambda}{d\nu}(t),z-v(t)\big\rangle\geq
0.
\end{equation*}
Indeed, let $t\in I$. By $(H_2)$, using Lemma \ref{lem2.4}, we can
ensure the existence of a sequence $(\zeta_n)_n$, such that
\begin{equation}\label{18}
\zeta_n\in
D\big(A(\theta_n(t),u_n(\phi_n(t)))\big),\;\;\;\zeta_n\longrightarrow
z\;\;\textmd{and}\;\; A^0\big(\theta_n(t),u_n(\phi_n(t))\big)
\zeta_n\longrightarrow A^0(t,u(t))z.
\end{equation}
Since $A(\theta_n(t), u_n(\phi_n(t)))$ is monotone, using
\eqref{12}, we have for each $n\in \mathbb{N}$ and $t\in I\setminus
J_n$,
\begin{equation}\label{19}
\big\langle v_n(\theta_n(t))-\zeta_n,
\frac{dv_n}{d\nu}(t)+f(t,u_n(\phi_n(t)),v_n(\phi_n(t)))\frac{d\lambda}{d\nu}(t)+A^0(\theta_n(t),
u_n(\phi_n(t)))\zeta_n\big\rangle\leq 0.
\end{equation}
Next,  let $t\in I\setminus \big(J\cup J'\cup(\underset{n}{\bigcup}
J_n)\big)$. From \eqref{17}, we have
  \begin{equation*}
  \big\langle \frac{dv}{d\nu}(t), v(t)-z\big\rangle\leq\limsup_{n\to\infty}\big\langle \frac{dv_n}{d\nu}(t),
 v(t)-z\big\rangle,
  \end{equation*}
and since from  \eqref{11} and \eqref{19}, we have for each $n\in
\mathbb{N}$,
\begin{eqnarray*}
&&\Big\langle \frac{dv_n}{d\nu}(t),
  v(t)-z\Big\rangle\leq\Big\langle
  \frac{dv_n}{d\nu}(t),v(t)-v_n(\theta_n(t))\Big\rangle+ \Big\langle\frac{dv_n}{d\nu}(t), v_n(\theta_n(t))-\zeta_n\Big\rangle+\Big\langle
  \frac{dv_n}{d\nu}(t),\zeta_n-z\Big\rangle\\&\leq&
  M_5\big(\|v(t)-v_n(\theta_n(t))\|+\|\zeta_n-z\|\big)+\Big\langle
  \frac{dv_n}{d\nu}(t),v_n(\theta_n(t))-\zeta_n\Big\rangle
  \\&\leq&
  M_5\big(\|v(t)-v_n(\theta_n(t))\|+\|\zeta_n-z\|\big)\\&+&\Big\langle f\big(t,u_n(\phi_n(t)), v_n(\phi_n(t))\big)\frac{d\lambda}{d\nu}(t)+A^0(\theta_n(t),
u_n(\phi_n(t)))\zeta_n,\zeta_n-v_n(\theta_n(t))\Big\rangle
\end{eqnarray*}
we conclude, using \eqref{eq4'}, \eqref{lip} and \eqref{18}
that\begin{equation} \big\langle \frac{dv}{d\nu}(t),
v(t)-z\big\rangle\leq\big\langle
f(t,u(t),v(t))\frac{d\lambda}{d\nu}(t)+A^0(t,
u(t))z,z-v(t)\big\rangle,
\end{equation}
which, with the fact that $v(t)\in D(A(t,u(t)))$, $u(0)=u_0$,
$v(0)=v_0$ and \eqref{noter}, means that $(u,v)$ is a BVRC solution
to our problem $(P_f)$. $\hfill\square$ \vskip2mm

\begin{rem}
If the operators depend only on the time and satisfy the following
hypotheses:\\
$(H'_1)$ There exists a function $\rho:I\longrightarrow [0,
+\infty[$, which is right-continuous on $[0, T[$ and nondecreasing
with $\rho(0) = 0$ and $\rho(T)<+\infty$ such that
\begin{equation*}dis(A(t), A(s))\leq d\rho(]s,
t])\;\;\textmd{for}\;\;0\leq s\leq t\leq T.\end{equation*}
 $(H'_2)$ There
exists a nonnegative real constant $c$ such that
\begin{equation*}\|A^0(t)x\|\leq c(1+\| x\|)\;\;\textmd{for}\;\;t\in
I,\;x\in D(A(t)).\end{equation*}
 $(H'_3)$ For any $t\in I$, $D(A(t))$
 is relatively ball-compact.\\
 Our theorem reads as
\end{rem}

\begin{coro}\label{Theorem 3.2}    Let for
every $t\in I$, $A(t):D(A(t))\subset \mathcal{H}\rightrightarrows
\mathcal{H}$ be a maximal monotone operator satisfying $(H'_1)$,
$(H'_2)$ and $(H'_3)$. Let $f:I\times \mathcal{H}\times
\mathcal{H}\longrightarrow \mathcal{H}$  such that $(H_4)$, $(H_5)$
and $(H_6)$ are satisfied. Then for any $(u_0,v_0)\in
\mathcal{H}\times D(A(0))$, there exists a BVRC solution $ (u, v) :I
\longrightarrow \mathcal{H}\times \mathcal{H}$ to  the problem
$$
\begin{cases}
u(0)=u_0, v(0)=v_0\in D(A(0));\\
 v(t)\in D(A(t))\;\;\;\forall t\in I;\\
\displaystyle\frac{du}{d\nu }(t)=v(t)\;\;\;d\nu-a.e.\,t\in
   I;\\
   -\displaystyle\frac{dv}{d\nu }(t)\in A(t) v(t)+ f(t,u(t), v(t))  \frac{d\lambda} {d\nu}(t)  \;\;\;d\nu-a.e.\,t\in
   I.
\end{cases}
$$
\end{coro}
\begin{rem}
In particular, if we consider $A(t,x)=N_{C(t,x)}$, the normal cone,
in the sense of convex analysis, to the closed convex set $C(t,x)$,
then under the following hypotheses:\\ $(H''_1)$ There exists a
function $\rho:I\longrightarrow [0, +\infty[$, which is
right-continuous on $[0, T[$ and nondecreasing with $\rho(0) = 0$
and $\rho(T)<+\infty$ and a nonnegative real constant $\gamma$ such
that
\begin{equation*}
|d_{C(t,x)}(z)-d_{C(s,y)}(z)|\leq d\rho(]s,
t])+\gamma\|x-y\|\frac{d\lambda}{d\nu}(s)\;\;\textmd{for}\;\;0\leq
s\leq t\leq T,\;x,y,z\in \mathcal{H}.
\end{equation*}
$(H''_2)$ For any bounded subset $E$ of $\mathcal{H}$, $C(I\times
E)$
 is relatively ball-compact.\\
 We can derive the following corollary.
\end{rem}
\begin{coro}
Let $C:I\times \mathcal{H}\rightrightarrows \mathcal{H}$ be a
set-valued map with nonempty, closed and convex values satisfying
$(H''_1)$ and $(H''_2)$. Let $f:I\times \mathcal{H}\times
\mathcal{H}\longrightarrow \mathcal{H}$  such that $(H_4)$, $(H_5)$
and $(H_6)$ are satisfied. Then for any $(u_0,v_0)\in
\mathcal{H}\times C(0,u_0)$, there exists a BVRC solution $ (u, v)
:I \longrightarrow \mathcal{H}\times \mathcal{H}$ to  the problem
$$
\begin{cases}
u(0)=u_0, v(0)=v_0;\\
 v(t)\in C(t, u(t))\;\;\;\forall t\in I;\\
\displaystyle\frac{du}{d\nu }(t)=v(t)\;\;\;d\nu-a.e.\,t\in
   I;\\
   -\displaystyle\frac{dv}{d\nu }(t)\in N_{C(t,u(t))} (v(t))+ f(t,u(t), v(t))  \frac{d\lambda} {d\nu}(t)  \;\;\;d\nu-a.e.\,t\in
   I.
\end{cases}
$$
\end{coro}

\end{document}